OpenStudy (anonymous):

Select the equation of a line that is perpendicular to the line on the graph and passes through the point (3,2)

OpenStudy (anonymous):

A. Y=1/3 x+3 B. Y=-1/3 x +3 C. Y=3x+2 D. Y=-1/3x+2

jimthompson5910 (jim_thompson5910):

please post the graph

OpenStudy (anonymous):

I posted an image with the question:)

OpenStudy (anonymous):

Here it is:)

jimthompson5910 (jim_thompson5910):

thanks

jimthompson5910 (jim_thompson5910):

what two points lie on this line?

OpenStudy (anonymous):

(0,-2)? Is that one

OpenStudy (anonymous):

Sorry I meant positive 2

jimthompson5910 (jim_thompson5910):

no, (0,-2) is not a point on the graph you posted

jimthompson5910 (jim_thompson5910):

yes (0,2) is

jimthompson5910 (jim_thompson5910):

what else?

OpenStudy (anonymous):

Is (1,5) one? Sorry I'm a little confused

OpenStudy (anonymous):

And (-1,-1)?

jimthompson5910 (jim_thompson5910):

(1,5) is another

jimthompson5910 (jim_thompson5910):

we only need 2 points to find the slope

jimthompson5910 (jim_thompson5910):

do you know the slope formula?

OpenStudy (anonymous):

Ok!:) what next?

jimthompson5910 (jim_thompson5910):

does this formula look familiar? \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (anonymous):

The slope is 3, correct?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Ok great!:) now what?

jimthompson5910 (jim_thompson5910):

think of 3 as 3/1 flip the fraction: 3/1 -----> 1/3 then flip the sign: +1/3 ----> -1/3 --------------------------------- the original slope is 3/1 the perpendicular slope is -1/3

jimthompson5910 (jim_thompson5910):

the perpendicular line has a slope of -1/3 and it goes through (3,2)

OpenStudy (anonymous):

Oh ok that makes sense, so it's either b or d?

jimthompson5910 (jim_thompson5910):

start with y = mx+b